Tohban Report 2012-07-25

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 18 July 2012
End Date: 25 July 2012
Tohban: Hazel Bain
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity has decreased this week due to AR 11520 rotating off the disk. However before it rotated out of view it produced one particularly nice M-class flare on the 19th of July. The solar background has now returned to low B-class level. There are currently 7 active regions on disk with STEREO B showing a couple of small active regions due to rotate onto the disk in the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      9    13     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      5    11     1     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      2     2     0     0

There were RHESSI flares/GOES flares 97 / 23 over the time range 18-Jul-12 25-Jul-12

Memory Management

SSR is currently ~16%. Increased low energy counts from detector 8 on exiting the SAA are thought to have contributed to a higher than normal SSR fill level for the current solar activity. The HV on detector 8 was was lowered by 5 steps (193 to 188) at 206/20:25 UT to try and resolve this problem. An extra one day of AGO passes have been scheduled for July 26th to help empty the SSR.

Spacecraft Status

The cryocooler power was decreased from 78 W to 71 W at 2012-200-21:03:56 UTC. The accelerations dropped from ~17 mG down to ~12 mG and the cold tip temperatures rose by a few tenths, during the remainder of the pass. The cold plate 1 and 2 temperatures are currently at 107.3 K and 105.1 K respectively.

A spin up was completed at 203/15:40 UT.

Data Gaps

There were no significant data gaps this week (only two ~5 min gaps on the 20th and 21st).

Detector issues

Due to an excess of low energy counts coming out of the SAA, the HV on detector 8 was lowered by 5 steps, from 193 to 188 at 206/20:25 UT. The data following this change has not yet been downlinked to check whether this has resolved the problem.

Other notes

Spacecraft Management

Decimation Active/vigorous (as of day 184)
HLAT Decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) +/- 4 minutes
Require extra passes? Yes
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detector 8 HV lowered by 5 steps (see above for details).
Personal tools